Dr. Paul Coddington,

Northeast Parallel Architectures Center,

111 College Place, 3-228 CST,

Syracuse University, Syracuse, NY 13244-4100.

Tel: (315) 443-4883 email: paulc@npac.syr.edu

11/1/93

To whom it may concern,

Chris Martin worked for myself and Professor James Allwright in a Research Experience for Undergraduates program in computational physics at Syracuse University in the summer of 1993. His project was to take some sequential graph coloring algorithms and programs and to rewrite them in two different parallel languages – Express Fortran (a MIMD message-passing language) and CM Fortran (a SIMD data-parallel language) – and to test their performance on two different parallel computers (an Intel iPSC/860 and a Thinking Machines CM-5) for some sample graphs.

This required an understanding of the graph coloring algorithms, the exisiting code, parallel computing concepts, and the two specific parallel languages and computers. In a very short period of time, Chris was able to familiarize himself with the algorithms and the parallel programming aspects of the work, and to produce working programs and results with very little supervision. He produced an impressive paper and poster on his work, and he will be co-author on a paper on this work to be submitted to an academic journal.

Chris is very quick to pick up new concepts, is a responsible and enthusiastic worker, and a good programmer. He is intelligent, amiable, polite, and a pleasure to work with. I believe he would do very well in graduate school and would recommend him highly.

Yours sincerely,